# The Limited Impact of Medical Adaptation of Large Language and Vision-Language Models

### v1

@@ 4.2.1 Selecting the “Right” Prompt Independently is Crucial for a Fair Comparison Between Models
In Figure 6, we show how the win/tie/loss rates change when we optimize the prompt only for the medical model, using Clinical-Camel-70B, OpenBioLLM-8B, and LLaVA-Med-v0-7B as an illustrative example. For Clinical-Camel-70B and OpenBioLLM-8B, we show the 3-shot results on the textual medical knowledge QA datasets (see Section 4.2.3 for discussion of results on textual clinical note QA, which are similar). For LLaVA-Med-v0-7B, we show the 3-shot results on all visual medical QA datasets. We note that the results here are shown with statistical testing, i.e., the win/tie/loss rates are calculated based on the 95% bootstrapping confidence intervals, as described in Section 3.
In all three cases, we observe that when we optimize the prompt only for the medical model, the win rates increase significantly and the loss rates decrease significantly. For example, for Clinical-Camel-70B (Figure 6(a)), we find that win rate increases from 0% to 54.5% and that the loss rate decreases from 36.4% to 0%, completely reversing the conclusion about Clinical-Camel-70B. We also observe that such trends exist regardless of model scale: across both 70B-parameter and 7/8B-parameter models, the results are highly sensitive to the choice of prompt format and few-shot examples.
Notably, we find that such differences in performance can arise as a result of seemingly insignificant differences in the prompt. As an illustrative example, we show the prompt formats optimized for Clinical-Camel-70B and its base model Llama-2-70B on MMLU-Medical (Clinical Knowledge):
@@
[frametitle=Prompt Formats for Clinical-Camel-70B vs. Llama-2-70B on MMLU-Medical (Clinical Knowledge)] Prompt Format Optimized for Clinical-Camel-70B:
@@
### Answer: We find that the main difference in the prompt, modulo the selected few-shot examples, lies on how the question and answer headers are formatted, with the overall semantics of the overall prompt remaining completely intact. We also note that the task instruction provided to Clinical-Camel-70B and Llama-2-70B is also identical: “The following is a multiple-choice question about medical knowledge. Answer the question by choosing one of the options from A to D.” These results are consistent with the conclusions of prior works that demonstrate that the behavior of LLMs/VLMs are highly sensitive to the specifics of the input prompt (Jiang et al., 2020; Zhao et al., 2021; Sclar et al., 2024), including works specifically focused on clinical tasks (Ceballos-Arroyo et al., 2024b).
Our findings suggest that for any pair of models, selecting the “right” prompt format and few-shot examples for each model separately is crucial to ensure a fair comparison in the zero-/few-shot prompting regime.
@@
Figure 7: Ignoring statistical uncertainty in model comparison can lead to overly optimistic or pessimistic conclusions about the performance benefits from medical DAPT: (a) Clinical-Camel-70B, (b) OpenBioLLM-8B, (c) LLaVA-Med-v0-7B. Here, we show the results when we separately perform model-specific prompt selection for each model in a given pair (“Independently Optimized”). “No Statistical Testing” refers to the setting where we compare models based on raw absolute accuracy. “Statistical Testing” refers to the setting where we compare models based on the 95% confidence intervals in relative accuracy. For Clinical-Camel-70B and OpenBioLLM-8B, we show the 3-shot win/tie/loss rates (%) on textual medical knowledge QA tasks (trends are similar for the textual clinical note QA tasks; see Section 4.2.3). For LLaVA-Med-v0-7B, we show the 3-shot win/tie/loss rates (%) on all visual QA tasks. Without statistical testing, both win and loss rates can be overestimated due to the failure to account for statistical ties.
In Figure 7, we show how the win/tie/loss rates change when we remove statistical testing, comparing the performances of each model pair solely based on raw absolute accuracy. As in Section 4.2.1, we show the 3-shot results for Clinical-Camel-70B and OpenBioLLM-8B on the textual medical knowledge QA datasets, and the 3-shot results for LLaVA-Med-v0-7B on all visual QA datasets. We note that the results here are shown after performing model-specific prompt selection for each model independently.
In all three cases, we overall observe that due to the absence of ties when comparing based on raw absolute accuracy, both the win and loss rates can become overestimated. In particular, for LLaVA-Med-v0-7B (Figure 7(c)), we observe that the win rate in the 3-shot setting increases from 12.5% to 62.5%, reversing the conclusions about its improvement over its base model LLaVA-v0-7B. For Clinical-Camel-70B and OpenBioLLM-8B, we find that while both models lose more than they win against their base models with or without statistical testing, the failure to account for statistical ties also leads to an overly pessimistic conclusion. In fact, with statistical testing, both models reach a tie in 60%+ of QA datasets when we account for statistical uncertainty (Figure 7(a) and (b)), which is almost entirely ignored when comparing based on raw accuracy. Accounting for statistical uncertainty allows one to more accurately conclude that all three models reach a statistical tie in most cases, with little to no statistically significant improvements in performance.
@@
Figure 8: Optimizing the prompt for only the medical model and comparing models without accounting for statistical uncertainty can overestimate the performance improvements from medical DAPT. We show the win/tie/loss rates (%) of the medical models across all (model pair, QA dataset) combinations, when (a) independently optimizing the prompt for each model and performing statistical testing, (b) optimizing the prompt only for the medical model and performing statistical testing, (c) independently optimizing the prompt for each model without statistical testing, and (d) optimizing the prompt only for the medical model without statistical testing. Here, we show the results for greedy decoding. The results for constrained decoding are similar (Figure E4).
@@
In Figure 8, we show how the win/tie/loss rates of the medical models, computed over all (model pair, QA dataset) combinations, vary across the four different settings. Overall, we find that for both LLMs and VLMs, the performance improvement from using a medically adapted model instead of its general-domain counterpart can be substantially overestimated when (i) the prompt is only tailored to the medical model; and (ii) the models are compared only based on their absolute accuracies. For example, we observe that the zero-shot win rate substantially increases from 9.4% to 70.5% on textual medical knowledge QA, 28.3% to 56.0% on textual clinical note QA, and 6.3% to 62.5% on visual QA, when only performing prompt selection for the medical model and comparing based on raw absolute accuracy. Figure E4 in Appendix E shows a similar trend in the win/tie/loss rates, when the model predictions are generated via constrained decoding.
In summary, these results highlight the importance of accounting for LLM/VLM sensitivity to the prompting details (Section 4.2.1) and statistical uncertainty (Section 4.2.2), in order to draw reliable conclusions about the effectiveness of medical DAPT.
5 Supervised Fine-Tuning Evaluation Results
Here, we summarize the main findings from the supervised fine-tuning experiments outlined in Section 3.2, focusing on the 7/8B-parameter medical LLMs and all medical VLMs. Overall, we find that after supervised fine-tuning, medical LLMs do show statistically significant improvements on the textual medical knowledge QA datasets (Section 5.1) but not on the textual clinical note QA datasets (Section 5.2), while medical VLMs do not show consistent improvements on visual QA (Section 5.3).
Table 5: The win, tie, and loss rates (%) of all medical LLMs on textual medical knowledge QA (left) and textual clinical note QA (right), after fine-tuning them on the training set of each QA dataset. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. We exclude the results for Med42-v2-8B on the textual medical knowledge QA datasets, as it has already been trained on all of them during medical DAPT.
@@ 5 Supervised Fine-Tuning Evaluation Results
Med42-v2-8B (Christophe et al., 2024b) - - - 12.5 75.0 12.5
OpenBioLLM-8B (Pal and Sankarasubbu, 2024) 25.0 75.0 0 0 100.0 0
MediTron-7B (Chen et al., 2023) 75.0 25.0 0 0 100.0 0
BioMistral-7B (Labrak et al., 2024) 50.0 50.0 0 0 100.0 0
BioMedGPT-LM-7B (Luo et al., 2023) 75.0 25.0 0 0 87.5 12.5
Aggregate 56.3 43.8 0 2.5 92.5 5.0
@@ 5.1 Evaluation of Medical LLMs on Textual Medical Knowledge QA
In Figure 9(a), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual medical knowledge QA datasets, after fine-tuning each model. In Table 5 (left), we also show the win, tie, and loss rates (%) of the medical LLMs on these datasets. We boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. As discussed in Section 3.2, we do not evaluate on the MMLU-Medical datasets, which only have 5 examples in the training set per subject (see Table A1). We also exclude the results for Med42-v2-8B, since it has already been trained on all medical knowledge QA datasets as part of medical DAPT (see Table 1).
Figure 9: Absolute and relative exact-match accuracies of (a) 7/8B-parameter LLMs on textual medical knowledge QA, (b) all VLMs on visual medical QA, and (c) 7/8B-parameter LLMs on clinical note QA, after supervised fine-tuning on each QA dataset. Overall, medical LLMs do show improvement in supervised fine-tuning performance on the textual medical knowledge QA datasets, but not on the textual clinical note QA datasets. Medical VLMs show some improvements on the visual QA datasets, which, however, are statistically insignificant.
Overall, we find that medical LLMs do consistently improve over their general-domain base models in the supervised fine-tuning regime. The top row of Figure 9(b) shows that all medical LLMs generally perform better than their base models in absolute terms. The bottom row of Figure 9(b) shows that in most cases, these improvements are indeed statistically significant, with the 95% confidence intervals lying above zero. In fact, Table 5 (left) shows that all medical LLMs outperform their base models with loss rates of 0%, winning much much more than they lose. Even BioMedGPT-LM-7B (brown), which significantly underperforms its base model Llama-2-7B in the zero-/few-shot prompting regime (see Sections 4.14.2), achieves a win rate of 75.0%. We also observe that medical models based on more recent general-domain models—OpenBioLLM-8B and BioMistral-7B, based on Llama-3-8B—tend to show lower win rates than the others based on older general-domain models (Llama-2-7B).
@@ 5.2 Evaluation of Medical LLMs on Textual Clinical Note QA
In Figure 9(c), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual clinical note QA datasets, after fine-tuning each model. In Table 5 (right), we also show the win, tie, and loss rates (%) of the medical LLMs on these datasets. We boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. As in the zero-/few-shot prompting evaluations in Section 4.1.2, we exclude the results on these datasets for MediTron-7B, as its small context window size of 2k tokens in insufficient for handling full-length clinical notes.
The top and bottom rows of Figure 9(c) clearly show that across all model pairs and datasets, the fine-tuning performances of the medical LLMs and their general-domain counterpart are almost exactly identical. In Table 5 (right), we find that across all (model pair, QA dataset) combinations, 92.5% of cases result in a statistical tie, with the medical models losing (5%) more than they win (2.5%) against their base models. In fact, we find that 4 out of 5 medical LLMs are completely indistinguishable from their base models in terms of the win/tie/loss rates, with the remaining model (BioMedGPT-LM-7B) losing more than it wins. Meanwhile, we observe that while the 7/8B-parameter LLMs do not consistently perform well off-the-shelf on clinical note QA (Section 4.1.2), both medical and general-domain models uniformly show strong performance after being fine-tuned towards each task.
@@ 5.3 Evaluation of Medical VLMs on Visual Medical QA
In Figure 9(b), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain VLMs on the visual QA datasets, after fine-tuning each model. In Table 6, we also show the win, tie, and loss rates (%) of the medical VLMs on these datasets. We boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. As noted in Section 3.2, we do not evaluate on the MMMU-Medical datasets, which only have 5 examples in the training set per subject (see Table A1).
Overall, we find that medical VLMs do not consistently improve over their general-domain base models in the supervised fine-tuning regime. In Figure 9(b), we find that LLaVA-Med-v0-7B (yellow) shows no statistically significant improvements over its base model, while Med-Flamingo-9B (cyan) shows improvement on PathVQA (+3.5% in raw accuracy) but reaches a tie in others. When we aggregate the results across all (model pair, QA dataset) combinations, the medical VLMs achieve win/tie/loss rates of 16.7%/66.7%/16.7% (Table 6), again indicating limited collective improvement. Meanwhile, given that these conclusions are based on a relatively small number of models (only 2) and datasets (only 3), we note that the generalizability of this finding may be limited.
Table 6: The win, tie, and loss rates (%) of all medical VLMs on visual QA, after fine-tuning them on the training set of each QA dataset. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa.
Model Visual QA
Win Tie Loss
LLaVA-Med-v0-7B (Li et al., 2023a) 0 66.7 33.3
Med-Flamingo-9B (Moor et al., 2023b) 33.3 66.7 0
Aggregate 16.7 66.7 16.7
In summary, the results in Sections 5.15.3 show that in the supervised fine-tuning regime, medical LLMs show improvements on textual QA tasks focused on assessing medical knowledge but not on those based on clinical notes, while medical VLMs generally fail to show statistically significant improvements on the visual medical QA datasets.

### v2

@@ 4.2.1 Selecting the “Right” Prompt Independently is Crucial for a Fair Comparison Between Models
In Figure 8, we show how the win, tie, and loss rates (%) change when we optimize the prompt only for the medical model, using Clinical-Camel-70B, OpenBioLLM-8B, and LLaVA-Med-7B as illustrative examples. For Clinical-Camel-70B and OpenBioLLM-8B, we show the 3-shot results on the textual medical knowledge QA datasets (see Section 4.2.3 for discussion of results on textual clinical note QA, which are similar). For LLaVA-Med-7B, we show the 3-shot results on all visual medical QA datasets. We note that the results here are shown with statistical testing, i.e., the win/tie/loss rates are calculated based on the 95% bootstrapping confidence intervals, as described in Section 3.
In all three cases, we observe that when we optimize the prompt only for the medical model, the win rates increase significantly and the loss rates decrease significantly. For example, we find that the win rate for Clinical-Camel-70B increases from 0% to 54.5% while its loss rate decreases from 36.4% to 0% under this setting, completely reversing the conclusion about Clinical-Camel-70B (Figure 8(a)). We also observe that such trends exist regardless of model scale: across both 70B-parameter and 7/8B-parameter models, the results are highly sensitive to the choice of prompt format and few-shot examples.
Notably, we find that such differences in performance can result from seemingly insignificant differences in the prompt. Below, we show the prompt formats optimized for Clinical-Camel-70B and Llama-2-70B on MMLU (Clinical Knowledge):
@@
[frametitle=Clinical-Camel-70B vs. Llama-2-70B on MMLU (Clinical Knowledge)] Prompt Format Optimized for Clinical-Camel-70B:
@@
### Answer: We find that the main difference in the prompt, modulo the selected few-shot examples, lies on how the question and answer headers are formatted, with the overall semantics of the overall prompt remaining completely intact. We also note that the task instruction provided to Clinical-Camel-70B and Llama-2-70B is also identical: “The following is a multiple-choice question about medical knowledge. Answer the question by choosing one of the options from A to D.” These results are consistent with the conclusions of prior works that demonstrate that the behavior of LLMs/VLMs are highly sensitive to the specifics of the input prompt (Jiang et al., 2020; Zhao et al., 2021; Sclar et al., 2024), including works specifically focused on clinical tasks (Ceballos-Arroyo et al., 2024).
Our findings suggest that selecting the “right” prompt format and few-shot examples for each model separately is crucial for a fair comparison in the zero-/few-shot prompting regime.
@@
Figure 9: Ignoring statistical uncertainty in model comparison can lead to overly optimistic or pessimistic conclusions about the performance benefits from medical DAPT: (a) Clinical-Camel-70B, (b) OpenBioLLM-8B, (c) LLaVA-Med-7B. Here, we show the results when we separately perform model-specific prompt selection for each model in a given pair (“Independently Optimized”). “No Statistical Testing” refers to the setting where we compare models based on raw absolute accuracy. “Statistical Testing” refers to the setting where we compare models based on the 95% confidence intervals in relative accuracy. For Clinical-Camel-70B and OpenBioLLM-8B, we show the 3-shot win/tie/loss rates (%) on the textual medical knowledge QA tasks (trends are similar for the textual clinical note QA tasks; see Section 4.2.3). For LLaVA-Med-7B, we show the 3-shot win/tie/loss rates (%) on all visual medical QA tasks. Without statistical testing, both win and loss rates can be overestimated due to the failure to account for statistical ties.
In Figure 9, we show how the win, tie, loss rates (%) change when we remove statistical testing, comparing the performances of each model pair solely based on raw absolute accuracy. As in Section 4.2.1, we show the 3-shot results for Clinical-Camel-70B and OpenBioLLM-8B on the textual medical knowledge QA datasets, and the 3-shot results for LLaVA-Med-7B on all visual medical QA datasets. We note that the results here are shown after performing model-specific prompt selection for each model independently.
In all three cases, we overall observe that due to the absence of ties when comparing based on raw absolute accuracy, both the win and loss rates can become overestimated. In particular, we observe that the win rate for LLaVA-Med-7B in the 3-shot setting increases from 12.5% to 62.5%, reversing the conclusions about its improvement over its base model LLaVA-v0-7B (Figure 9(c)). For Clinical-Camel-70B and OpenBioLLM-8B, we find that while both models lose more than they win against their base models with or without statistical testing, the failure to account for statistical ties also leads to an overly pessimistic conclusion. In fact, with statistical testing, both models reach a tie in 60+of QA datasets when we account for statistical uncertainty (Figures 9(a) and (b)), which is almost entirely ignored when comparing based on raw accuracy. Accounting for statistical uncertainty allows one to more accurately conclude that all three models reach a statistical tie in most cases, with little to no statistically significant improvements in performance.
@@
Figure 10: Optimizing the prompt for only the medical model and comparing models without accounting for statistical uncertainty can overestimate the performance improvements from medical DAPT. We show the win/tie/loss rates (%) of the medical models across all (model pair, QA dataset) combinations, when (a) optimizing the prompt for each model, with statistical testing; (b) optimizing the prompt only for the medical model, with statistical testing; (c) optimizing the prompt for each model, without statistical testing; and (d) optimizing the prompt only for the medical model, without statistical testing. We show the results for greedy decoding. The results for constrained decoding are similar (Figure D5).
@@
When we aggregate the results over all (model pair, QA dataset) combinations, we find that for both LLMs and VLMs, the performance improvement from medical DAPT can be substantially overestimated when (i) the prompt is only tailored to the medical model; and (ii) the models are compared only based on their absolute accuracies (Figure 10). For example, the aggregate zero-shot win rate substantially increases from 9.0% to 74.2% on textual medical knowledge QA, 28.0% to 60.0% on textual clinical note QA, and 6.3% to 62.5% on visual medical QA, when only performing prompt selection for the medical model and comparing based on raw absolute accuracy. We see a similar trend in the win/tie/loss rates when the model predictions are generated via constrained decoding (Figure D5).
In summary, these results highlight the importance of accounting for LLM/VLM sensitivity to the prompting details (Section 4.2.1) and statistical uncertainty (Section 4.2.2), in order to draw reliable conclusions about the effectiveness of medical DAPT (Section 4.2.3).
5 Supervised Fine-Tuning (SFT) Evaluation Results
We summarize the main findings from the SFT experiments outlined in Section 3.2. Overall, we find that after SFT, medical LLMs do show statistically significant improvements on the textual medical knowledge QA datasets (Section 5.1) but not on the clinical note QA datasets (Section 5.2), while medical VLMs do not show consistent improvements on the visual medical QA datasets (Section 5.3).
Table 6: The win, tie, and loss rates (%) of all medical LLMs on textual medical knowledge QA (left) and textual clinical note QA (right), after fine-tuning them on the training set of each QA dataset. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa. We exclude the results for the Med42 models on the textual medical knowledge QA datasets, as they have already been trained on all of those datasets during medical DAPT.
@@ 5 Supervised Fine-Tuning (SFT) Evaluation Results
Med42-v2-70B - - - 37.5 62.5 0
OpenBioLLM-70B 50.0 50.0 0 12.5 87.5 0
Med42-v1-70B - - - 0 100.0 0
MediTron-70B 75.0 25.0 0 0 100.0 0
Clinical-Camel-70B 0 100.0 0 0 100.0 0
Med42-v2-8B - - - 12.5 75.0 12.5
OpenBioLLM-8B 25.0 75.0 0 0 100.0 0
MediTron-7B 75.0 25.0 0 0 100.0 0
BioMistral-7B 50.0 50.0 0 0 100.0 0
BioMedGPT-LM-7B 75.0 25.0 0 0 87.5 12.5
Aggregate 53.8 46.2 0 6.7 90.7 2.7
@@ 5.1 Evaluation of Medical LLMs on Textual Medical Knowledge QA
In Figure 11(a), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual medical knowledge QA datasets, after fine-tuning each model. In Table 6 (left), we also show the win, tie, and loss rates (%) of the medical LLMs on these datasets. As discussed in Section 3.2, we do not evaluate on the MMLU-Medical datasets, which only have 5 examples in the training set per subject (Table A1). We also exclude the results for the Med42 models, since they have already been trained on all medical knowledge QA datasets via medical DAPT (Table 2).
Figure 11: Absolute and relative accuracies of (a) all LLMs on textual medical knowledge QA, (b) all VLMs on visual medical QA, and (c) all LLMs on textual clinical note QA, after SFT on each task. Overall, medical LLMs do show improvement in SFT performance on the textual medical knowledge QA datasets, but not on the textual clinical note QA datasets. Medical VLMs do not show significant improvements on the visual medical QA datasets.
Overall, we find that medical LLMs do consistently improve over their general-domain base models in the SFT regime. All medical LLMs generally perform better than their base models in absolute terms, albeit with marginal improvements (on average, 1.6% improvement in absolute accuracy) (top row of Figure 11(a)). In most cases, these improvements are nonetheless statistically significant, with the 95% confidence intervals in relative accuracy lying above zero (bottom row of Figure 11(a)). In fact, all medical LLMs achieve loss rates of 0%, with 5 out of 7 models achieving win rates of 50+(Table 6, left). Even BioMedGPT-LM-7B (brown), which significantly underperforms its base model Llama-2-7B-Chat in the zero-/few-shot prompting regime (Sections 4.14.2), achieves a win rate of 75.0%. We also find that medical models based on more recent general-domain base models (e.g., OpenBioLLM-70B and OpenBioLLM-8B, based on Llama-3) tend to show lower win rates than the other medical models based on older general-domain base models (e.g., MediTron and BioMedGPT-LM, based on Llama-2).
@@ 5.2 Evaluation of Medical LLMs on Textual Clinical Note QA
Figure 12: Even after accounting for the imbalance in the distribution of clinical acronyms in the CASI and MIMIC-III sense disambiguation datasets, medical LLMs fail to consistently improve over their base models, after SFT. In each panel, the top row shows the absolute macro exact-match accuracies and F1 scores—averaged over clinical acronyms—on the test set, and the bottom row shows the relative macro exact-match accuracies and F1 scores along with 95% confidence intervals derived via bootstrapping on the test set (Section 3).
In Figure 11(c), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain LLMs on the textual clinical note QA datasets, after fine-tuning each model. For the CASI and MIMIC-III sense disambiguation datasets, we also show the absolute and relative macro exact-match accuracies and F1 scores—averaged over clinical acronyms—in Figure 12 to account for the imbalance in the acronym distribution (Section 4.1.2). In Table 6 (right), we also show the win, tie, and loss rates (%) of the medical LLMs on these datasets. As in the zero-/few-shot prompting evaluations discussed in Section 4.1.2, we exclude the results for MediTron-7B (and its base model Llama-2-7B) on the EHRNoteQA and i2b2 datasets, as its small context window size of 2k tokens is insufficient for handling the full-length clinical notes in these datasets.
Across all model pairs and datasets, we find that the fine-tuning performances of the medical LLMs and those of their base models are nearly indistinguishable (Figure 11(c)). In Figure 12, we make similar observations on the CASI and MIMIC-III sense disambiguation datasets even after accounting for the imbalance in the clinical acronym distribution. In Table 6 (right), we find that across all (model pair, textual clinical note QA dataset) combinations, 90.7% of cases result in a statistical tie, with the medical models winning against their base models in only 6.7% of cases. In fact, 7 out of 10 medical LLMs are indistinguishable from their base models in terms of the win/tie/loss rates. Meanwhile, we observe that while the LLMs do not consistently perform well off-the-shelf on clinical note QA (Section 4.1.2), both medical and general-domain models uniformly show strong performance after being fine-tuned for each task, across all model scales.
@@
Table 7: The win, tie, and loss rates (%) of all medical VLMs on visual medical QA, after fine-tuning them on the training set of each QA dataset. For each medical model, we boldface the win rate if it wins more than it loses to its general-domain base model, and vice versa.
Model Visual Medical QA
Win Tie Loss
LLaVA-Med-7B 0 66.7 33.3
Med-Flamingo-9B 33.3 66.7 0
Aggregate 16.7 66.7 16.7
@@ 5.3 Evaluation of Medical VLMs on Visual Medical QA
In Figure 11(b), we show the absolute and relative exact-match accuracies achieved by the medical and general-domain VLMs on the visual medical QA datasets, after fine-tuning each model. In Table 7, we also show the win, tie, and loss rates (%) of the medical VLMs on these datasets. As discussed in Section 3.2, we do not evaluate on the MMMU-Medical datasets, which only have 5 examples in the training set per subject (Table A1).
Overall, medical VLMs do not consistently improve over their general-domain base models in the SFT regime. We observe that LLaVA-Med-7B (light green) shows no statistically significant improvements over its base model, while Med-Flamingo-9B (cyan) shows improvement on PathVQA (+3.5% in raw accuracy) but reaches a tie in others (Figure 11(b)). When we aggregate the results across all (model pair, visual medical QA dataset) combinations, the medical VLMs achieve win/tie/loss rates of 16.7%/66.7%/16.7% (Table 7), indicating limited collective improvement. Meanwhile, given that these conclusions are based on a relatively small number of models (only 2) and datasets (only 3), we note that the generalizability of this finding may be limited.
In summary, the results in Sections 5.15.3 show that in the SFT regime, medical LLMs show improvements on textual QA tasks focused on assessing medical knowledge but not on those based on clinical notes, while medical VLMs generally fail to show statistically significant improvements on the visual medical QA datasets.